Montserrat Volcano Observatory

Daily Report
Report for the period 6 pm 7 January
to 6 pm 8 January 1998

Activity at the volcano today was very similar to that of yesterday. Between 16:00 yesterday and 16:00 today there were 44 hybrid earthquakes, 38 rockfall signals, 15 long-period earthquakes and 6 volcano-tectonic earthquake. The 6 to 8 hour cyclicity of recent days appears to have continued today.

Cloud continues to mask the upper parts of the dome. However, rockfall activity on the southwestern flank of the volcano suggests continued rapid dome growth in the Galway's sector. Visual observations of the area where the Galway's Soufriere used to be show that new rock fall material is already filling up the depression left after the Boxing Day dome collapse and landslide. Activity on other sectors of the dome is at a low level.

Dr. Jenni Barclay left Montserrat after working as senior scientist over the Holiday period. Mr. Richie Robertson has arrived and takes over as Chief Scientist today.

Air quality at all monitoring stations across the island is good today.

People should remain alert and listen to Radio Montserrat for further information.

Montserrat Volcano Observatory